• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 08:04:30 on April28,2004

    Entry number 119675

    keyword=Owl Shift Summary

    Continue data taking with Be target
    01:10 - J.Segal called and asked to turn on gas flushing in the RICH.
    As a consequence HV VDC L tripped and was restored.
    01:40 - ROC12,ROC16 hung up - took about 15 min to restart
    02:50 - RICH: We noticed a few "red" numbers on the RICH gas panel:
    Argon flow (0), Radiator oxygen (-1.1ppm), radiator moisture
    (386ppm). We called Guido, he asked to call Jack, then
    Guido called and told that the it was no problem.
    04:40 - Beam is tripping, The current is reduced to 84uA.
    05:50-06:35 - Network switch in the hall tripped and had to
    be reset (Control Access).
    06:47 - data taking resumed

    Runs Taken



    Run time Comment
    2858 .....-00:30 1M events
    2959 00:30-00:52 1M events,
    2960 00:53-01:13 1M events. VDC L tripped for about 1minute.
    2961 01:14-01:37 1M events
    2962 - does not exist
    2963 01:53-02:13 1M events
    2964 02:14-02:37 1M events
    2965 02:37-02:58 1M events
    2966 02:58-03:21 1M events
    2967 03:23-03:43 1M events
    2968 03:44-04:09 1M events
    2969 04:09-04:30 1M events
    2970 04:31-04:52 380K events
    2971 04:55-05:29 1M events
    2972 05:30-05:40 340k events (stopped because of the network failure)
    2973 06:47-07:13 1M events
    2974 07:14-07:46 1M events


    Summary


    15.6M events have been taken with the "thick" Be target.